# Understanding the chemistry of temperate exoplanets atmospheres through experimental and numerical simulations

### v1

@@ 3.2.2 Oxidized organic compounds
Figure 7: Predicted abundance values with 0D-simulations of the oxidized organics in the different gas mixtures.
@@ 3.3 Formation mechanisms
From the 0D simulations performed with ReactorUI, we can determine the chemical pathways explaining the formation of the compounds of interest.
@@ 3.3.1 Reduced organic compounds
This chemical reactions network (Figure 8(a)) is consistent with the well-known methane chemistry and hydrocarbon formation studied in the reduced atmospheres of our solar system, such as Titan (Nixon_2024), Jupiter (Moses_2005) or Saturn (Moses_2000), as well as with predictions made for exoplanets with H2 dominant atmospheres (Adams_2022). This validates our approach, and gives us confidence for the study of our more oxidized atmospheric ranges.
Figure 8(b) shows the formation pathway of reduced organics from CO and CO2. Organic growth is inhibited compared to reduced mixtures, as has been previously studied in the literature — particularly in the CO2 case by Trainer_2004. It requires passing through more intermediates than in the reduced atmospheres, which limits the formation of reduced organic compounds. The formation of C2H4 and C2H6 requires the prior formation of the CH3 intermediate, which in turn is produced by two consecutive intermediates (C and 3CH2) after CO dissociation. This long reaction path explains the lower abundance of hydrocarbons compared to the reduced atmosphere. Furthermore, the dissociation of CO, which provides a source of carbon atoms for organic growth, is at the same time a source of exited oxygen radicals, mainly in the O(3P) state, which will oxidize some of the carbonate radical intermediates and reduced organics (Figure 9). For example, 3CH2 reacts with O(3P) to reform CO. However, as the concentration of CO in gas mixtures increases, there is an overall rise in the formation of reduced organic compounds, showing that with CO, the effect of increasing the carbon source dominates over the loss of intermediates by oxidation for the formation of the organics.
@@
As CH4 is a reactant in the CH4/CO/CO2 mixtures, the reaction pathway presented in Figure 8(a) also explains the organic growth. When atmospheres are more oxidized, we can observe the same mechanisms that are distinctive of fully oxidized atmospheres, as developed in Figure 8(b), with however a limitation in organic growth due in particular to the oxidation of reaction products and intermediates.
@@
Figure 8: Formation pathways of reduced organics with main (solid arrows) and minor (dotted arrows) reactions, started from a reduced reactant (CH4 8(a)) or an oxidized one (CO or CO2 8(b)).
@@ 3.3.2 Oxidized organic compounds
The main formation pathways of H2CO and CH3CHO are different depending on the oxidized reactant, as shown in Figure 10(a) and 10(c), while for CH3OH it is similar for CO and CO2. All these reaction paths involve reduced organic intermediates, such as CH3 and 3CH2, for which the formation efficiency in CO or CO2 mixtures is described in section 3.2.1. This explains the formation trends observed, i.e. formation efficiency is better with CO than with CO2, enhanced at high CO concentrations and limited at high CO2 concentrations. The formation of CH3OH and CH3CHO in mixtures of CO and CO2 requires at least one more intermediate than the formation of H2CO. This makes the formation of H2CO up to six orders of magnitude faster, explaining its predominance.
@@
$H_{2}CO+h\nu\longrightarrow CO+H_{2}$ (3)
The reaction of H2CO with H is almost two orders of magnitude faster than its photodissociation in the 0D model. HCO and CO are themselves involved as reactants in the formation of H2CO. In this way, H2CO is self-sustaining, which explains its high abundance. In contrast, CH3OH can be photo-dissociated through:
$CH_{3}OH+h\nu\longrightarrow H_{2}CO+H_{2}$ (4)
@@
Figure 10: Formation pathways of oxidized organics. The formation pathways of the intermediates CH3 and 3CH2 from CO or CO2 are detailed in Figure 8 and are then simplified here by long gray dotted arrows. The main (solid arrows) and minor (short dotted arrows) reactions are represented.
@@ 4 Discussion
4.1 Roles of methane, carbon monoxide and carbon dioxide in organic growth
@@
Methane plays a predominant role in organic growth, as demonstrated in our experiments, in which we observed a positive correlation between gas-phase hydrocarbon production and methane concentration. This is consistent with previous experiments in the literature, which have highlighted the major role of methane in the growth of solid particles (tholins) when N2/CH4 Titan-type mixtures are irradiated (Horst_2013; Horst_2018; Moran_2022; Jovanovic_2020; Jovanovic_2021). The production of the solid phase in Titan-type mixtures is optimized between 3 and 5% of CH4 (Sciamma_2010). In this study, in H2-dominated mixtures up to 10% of CH4, we observe continuous growth of hydrocarbons in the gas phase, without an obvious optimal concentration. This optimum would possibly occur at higher proportions of CH4.
@@
While methane is clearly identified as a major driver of organic growth, the effect of oxidized reactants is not as clearly correlated. At first glance, we identified an inhibitory effect due to the atomic oxygen source represented by the oxidized reactants, thus limiting the growth of long carbon chains. This is even more pronounced with CO2, which provides twice as much oxygen. This effect has also been identified in the production of solid particles in N2/CH4 mixtures, where the addition of CO alters the kinetics of methane, delaying the appearance of solid particles and reducing their production rate (Fleury_2014; He_2017; Moran_2022).
@@
However, CO and CO2 are also sources of carbon and can contribute to organic growth. Trainer_2006 showed that adding CO2 to a N2/CH4 mixture could increase particle production up to a CH4/CO2 ratio of 1 before decreasing. This effect is particularly evident in our CH4-free mixtures, where hydrocarbon production is still observed, although at lower levels than in the case where CH4 is present. The addition of CO to N2/CH4 mixtures, although reducing the production rate of tholins, increases their size (Fleury_2014; HorstTolbert_2014). In our study, we observe a positive correlation between the formation of organic compounds in the gas phase and the concentration of CO. That highlights that with CO, the effect of increasing the carbon source dominates over the loss by oxidation. In the case of a low methane concentration (1%), Jovanovic_2020 observed a higher production rate in the presence of 500 ppm CO than in its absence, highlighting the important role that CO can play as a carbon source when methane is less dominant.
@@
Furthermore, the addition of oxygen atoms increases the diversification and organic complexity of photochemical products. We observe the appearance of oxidized organic compounds of prebiotic interest in our CO and CO2-rich mixtures (CH3OH, H2CO, CH3CHO). It has also been shown that adding CO to Titan-type mixtures improves the reactivity of the gas phase (Yang_2025; Fleury_2014), leading to the incorporation of oxygen in both the gas and solid phases (He_2017). Similarly, the addition of CO2 leads to more oxidized solid compounds (Trainer_2004). Since CO2 is a more important source of oxygen than CO, one would expect oxygen incorporation to be better. In our study, this corresponds to the observation that oxidized organic compounds are better produced in the gas phase with CO2 than with CO when combined with CH4, and therefore when organic growth is not too limited. For the solid phase, Drant_2025 observed in N2/CH4-rich exoplanetary haze analogues that the characteristic absorption bands of oxygenated functions (O–H, C=O) are clearly visible in analogues produced with CO2, but not or only very weakly in those produced with CO alone.
@@ 4.2 Photochemistry of hydrocarbons in temperate methane-rich exoplanet atmosphere
Experiments conducted in reduced gas mixtures reveal similar formation rates for C2 compounds, preventing any definitive conclusions about selectivity. Simulated abundance using the 0D model also predicts comparable values for these molecules, but with a slight selectivity observed for C2H6 and C2H4 relative to C2H2. This contrasts with the hydrocarbon photochemistry observed in the gas giants Jupiter and Saturn, where detected abundance is depleted in C2H4 and enriched in C2H6 and C2H2 (Moses_2000; Moses_2005; Nixon_2010). These planets have hydrogen-dominated atmospheres with significant helium concentrations and a small fraction of methane (around 0.1%). They are cooler — around 150 K at 1 mbar.
The low abundance of C2H4 is explained by its high photodissociation rate at Lyman-$\alpha$, producing C2H2, and by its reaction with H, which recycles it into CH4 (Gladstone_1996). In contrast, C2H6 is shielded by methane and thus remains stable at Lyman-$\alpha$. In our study, the energy distribution of electrons in our plasma covers a much wider range than that of photons in planetary atmospheres, which are largely concentrated around Lyman-$\alpha$. Consequently, the wavelength selectivity of photodissociation is not reproduced, which likely contributes to the differences observed in the relative abundance of C2 compounds between our experiments and planetary atmospheres.
Additionally, C2H2, as a highly unsaturated compound, would normally be unstable in an H2-dominated environment. However, it is efficiently recycled via the recombination reaction:
$C_{2}H+H_{2}\longrightarrow C_{2}H_{2}+H$ (5)
We have noticed that this reaction is negligible in our reduced gas mixtures, whereas it becomes significant in oxidized mixtures. Since CO and CO2 are less reducing sources of carbon than CH4, this increases the degree of unsaturation of the products. This has also been demonstrated for the production of solid products (He_2017; Yang_2025). This corresponds to the fact that we observe a depletion of C2H2 in mixtures containing only CH4, whereas there is no apparent selectivity in mixtures containing CO and CO2.
Temperature differences may also play a role in C2 relative abundance. Adams_2022 suggests that warmer H2-rich atmospheres, compared to Titan, favor dynamic chemical cycles involving C2H2, C2H3, and C2H4 through reactions with atomic hydrogen. These processes can reduce C2H6 and C2H2 concentrations while increasing the relative abundance of C2H4.
@@ 4.3.1 Hydrocarbons
The C2 compounds appear to be relatively abundant and efficiently produced in our experiments, which are dominated by out-of-equilibrium chemistry. Similarly, a 1D non-equilibrium chemical model fitting the JWST observational data of K2-18 b predicts that C2H2, C2H4 and C2H6 could reach volume mixing ratios of up to a few percent in the upper atmosphere (at 1 mbar) (Jaziri_2025). C2H2 and C2H4 are in principle detectable with JWST, using combined observations from the NIRSpec/G395 and MIRI LRS instruments, provided their mixing ratios exceed 10-6 (Gasman_2022). This suggests that the detection of hydrocarbons in the atmosphere of K2-18 b is plausible if methane is abundant and predominant over carbon dioxide, as predicted by Wogan_2024; Schmidt_2025; Jaziri_2025; Hu_2025, and if photochemistry is sufficiently active. The 3.4 and 6.9 µm features observed in the NIRSpec and MIRI data could plausibly be attributed to C2H6 (Luque_2025), providing an alternative to their attribution to DMS (dimethyl sulfide) or DMDS (dimethyl disulfide) by Madhusudhan_2025. In summary, two-atom hydrocarbons are good candidates for detection with current JWST observations and future high-resolution ground-based observations, both in terms of expected abundance and photochemical plausibility. It is necessary to consider them before considering more complex species with a variety of potential sources.
@@ 4.3.2 Oxidized organics
Our complementary experimental approach, combining Mass Spectrometry and InfraRed spectroscopy, enables robust identification of H2CO, CH3OH, and CH3CHO formation in oxidized gas mixtures. Jaziri_2025 predict that H2CO, CH3OH, and CH3CHO could reach volume mixing ratios of up to a few ppm (10-6) in the upper atmosphere of K2-18 b. However, detectability studies with JWST indicate that species such as H2CO and CH3OH require volume mixing ratios exceeding 10-5 to produce measurable features in H2-dominated atmospheres (Huang_2022; Zhan_2022). Their detection in the atmosphere of K2-18 b therefore seems unlikely given current sensitivity limits. Nevertheless, our results suggest that in other H2-rich exoplanets with a higher CO and/or CO2 content, the formation of oxidized organic species may be particularly efficient. In such environments, these molecules could form more abundantly and potentially reach detectable levels with JWST or with future high-resolution ground-based instruments on the Extremely Large Telescope (ELT), such as ANDES (ArmazoNes high Dispersion Echelle Spectrograph, ANDES_2025) and PCS (Planetary Camera and Spectrograph, PCS_2021).
@@ 5 Conclusion
We studied out-of-equilibrium chemistry in H2-dominant atmospheres typical of temperate sub-Neptunes, combining experimental and numerical simulations. This approach complements previous laboratory studies by enabling more accurate molecular identification and clearer elucidation of reaction pathways. Our experiments, which are based on the combined use of mass spectrometry and infrared spectroscopy, have enabled the reliable detection of reduced (C2H2, C2H4, C2H6) and oxidized (CO, CO2, H2O, H2CO, CH3OH, CH3CHO) species in gas mixtures covering a wide range of C/O ratios and metallicities. The use of the 0D photochemical model allowed us to extract abundance trends and identify formation pathways.
Hydrocarbon compounds are produced, particularly in methane-rich atmospheres. The efficiency of this organic growth increases with metallicity in CH4-rich and CO-rich atmospheric analogues. In CO2-rich atmospheric analogues, oxidative loss reactions become predominant and limit the formation of hydrocarbons. The formation of hydrocarbons from methane can be explained by reaction pathways that have already been well studied in the atmospheres of the solar system, whereas their formation from oxidized reactants — CO or CO2 — involves more reaction intermediates and a change in the predominance of certain pathways. This explains the lower abundance of reduced organic compounds in an oxidized atmosphere, combined with the emergence of oxidation pathways that consume reaction intermediates, which have more influence for low C/O ratios. The addition of oxygen diversifies the chemistry and leads in particular to the formation of the oxidized organic compounds presented above, in CO-rich and CO2-rich mixtures. The combination of CH4 and CO is two-fold. On the one hand, it provides carbon-rich atmospheres, which promote the efficient formation of hydrocarbons. On the other hand, it leads to moderately oxygenated atmospheres, which promote the formation of oxidized organic compounds without overly enhancing the pathways for hydrocarbon destruction.
Our study highlights the importance of adopting a synergistic approach combining experiments, modeling, and observations in order to deepen our understanding of exoplanet atmospheres. Each field offers unique perspectives, but through close collaboration, their potential is greatly enhanced. It is essential to strengthen these interdisciplinary efforts.

### v2

@@ 3.2.2 Oxidized organic compounds
Figure 7: Predicted abundance values with 0D simulations of the oxidized organics in the different gas mixtures.
@@ 3.3 Formation mechanisms
From the 0D simulations performed with ReactorUI, we can determine the key reactions involved in the production and destruction of each chemical compound. This allows us to reconstruct the chemical networks involving the molecules that interest us in our study.
@@ 3.3.1 Reduced organic compounds
This chemical reactions network (Figure 8(a)) is consistent with the well-known methane chemistry and hydrocarbon formation studied in the reduced atmospheres of our Solar System, such as Titan (Nixon, 2024), Jupiter (Moses et al., 2005) or Saturn (Moses et al., 2000), as well as with predictions made for exoplanets with H2 dominant atmospheres (Adams et al., 2022). This validates our approach, and gives us confidence for the study of our more oxidized atmospheric ranges.
Figure 8(b) shows the formation pathway of reduced organics from CO and CO2. Organic growth is inhibited compared to reduced mixtures, as has been previously studied in the literature — particularly in the CO2 case by Trainer et al. (2004). It requires passing through more intermediates than in the reduced atmospheres, which limits the formation of reduced organic compounds. The formation of C2H4 and C2H6 requires the prior formation of the CH3 intermediate, which in turn is produced by two consecutive intermediates (C and 3CH2) after CO dissociation. This long reaction path explains the lower abundance of hydrocarbons compared to the reduced atmosphere. Furthermore, the dissociation of CO, which provides a source of carbon atoms for organic growth, is at the same time a source of exited oxygen radicals, mainly in the O(3P) state, which will oxidize some of the carbonate radical intermediates and reduced organics (Figure 9). For example, 3CH2 reacts with O(3P) to reform CO. However, as the concentration of CO in gas mixtures increases, there is an overall rise in the formation of reduced organic compounds, showing that with CO, the effect of increasing the carbon source dominates over the loss of intermediates by oxidation for the formation of the organics.
@@
As CH4 is a reactant in the CH4/CO/CO2 mixtures, the reaction pathway presented in Figure 8(a) also explains the organic growth. When atmospheres are more oxidized, we can observe the same mechanisms that are distinctive of fully oxidized atmospheres, as developed in Figure 8(b), but with a limitation in organic growth due to the oxidation of reaction products and intermediates.
@@
Figure 8: Formation pathways of reduced organics with main (solid arrows) and minor (dotted arrows) reactions started from a reduced reactant (CH4 8(a)) or an oxidized one (CO or CO2 8(b)).
@@ 3.3.2 Oxidized organic compounds
The main formation pathways of H2CO and CH3CHO are different depending on the oxidized reactant, as shown in Figure 10(a) and 10(c), while for CH3OH it is similar for CO and CO2. All these reaction paths involve reduced organic intermediates, such as CH3 and 3CH2, for which the formation efficiency in CO or CO2 mixtures is described in section 3.2.1. This explains the formation trends observed, i.e., formation efficiency is better with CO than with CO2, enhanced at high CO concentrations and limited at high CO2 concentrations. The formation of CH3OH and CH3CHO in mixtures of CO and CO2 requires at least one more intermediate than the formation of H2CO. This makes the formation of H2CO up to six orders of magnitude faster, explaining its predominance.
@@
$H_{2}CO+h\nu\longrightarrow CO+H_{2}.$ (3)
The reaction of H2CO with H is almost two orders of magnitude faster than its photodissociation in the 0D model. HCO and CO are themselves involved as reactants in the formation of H2CO. In this way, H2CO is self-sustaining, which explains its high abundance. In contrast, CH3OH can be photo-dissociated through
$CH_{3}OH+h\nu\longrightarrow H_{2}CO+H_{2},$ (4)
@@
Figure 10: Formation pathways of oxidized organics. The formation pathways of the intermediates CH3 and 3CH2 from CO or CO2 are detailed in Figure 8 and are simplified here by long dotted gray arrows. The main (solid arrows) and minor (short dotted arrows) reactions are represented.
@@ 4 Discussion
4.1 Roles of methane, carbon monoxide, and carbon dioxide in organic growth
@@
Methane plays a predominant role in organic growth, as demonstrated in our experiments, in which we observed a positive correlation between gas-phase hydrocarbon production and methane concentration. This is consistent with previous experiments in the literature, which have highlighted the major role of methane in the growth of solid particles (tholins) when N2/CH4 Titan-type mixtures are irradiated (Hörst and Tolbert, 2013; Hörst et al., 2018; Moran et al., 2022; Jovanović et al., 2020; Jovanovic, 2021). The production of the solid phase in Titan-type mixtures is optimized between 3 and 5% of CH4 (Sciamma-O’Brien et al., 2010). In this study, in H2-dominated mixtures up to 10% of CH4, we observe continuous growth of hydrocarbons in the gas phase, without an obvious optimal concentration. This optimum would possibly occur at higher proportions of CH4.
@@
While methane is clearly identified as a major driver of organic growth, the effect of oxidized reactants is not as clearly correlated. At first glance, we identified an inhibitory effect due to the atomic oxygen source represented by the oxidized reactants, thus limiting the growth of long carbon chains. This is even more pronounced with CO2, which provides twice as much oxygen. This effect has also been identified in the production of solid particles in N2/CH4 mixtures, where the addition of CO alters the kinetics of methane, delaying the appearance of solid particles and reducing their production rate (Fleury et al., 2014; He et al., 2017; Moran et al., 2022).
@@
However, CO and CO2 are also sources of carbon and can contribute to organic growth. Trainer et al. (2006) showed that adding CO2 to a N2/CH4 mixture could increase particle production up to a CH4/CO2 ratio of 1 before decreasing. This effect is particularly evident in our CH4-free mixtures, where hydrocarbon production is still observed, although at lower levels than in the case where CH4 is present. The addition of CO to N2/CH4 mixtures, although reducing the production rate of tholins, increases their size (Fleury et al., 2014; Hörst and Tolbert, 2014). In our study, we observe a positive correlation between the formation of organic compounds in the gas phase and the concentration of CO. That highlights that with CO, the effect of increasing the carbon source dominates over the loss by oxidation. In the case of a low methane concentration (1%), Jovanović et al. (2020) observed a higher production rate in the presence of 500 ppm CO than in its absence, highlighting the important role that CO can play as a carbon source when methane is less dominant.
@@
Furthermore, the addition of oxygen atoms increases the diversification and organic complexity of photochemical products. We observe the appearance of oxidized organic compounds of prebiotic interest in our CO and CO2-rich mixtures (CH3OH, H2CO, CH3CHO). It has also been shown that adding CO to Titan-type mixtures improves the reactivity of the gas phase (Yang et al., 2025; Fleury et al., 2014), leading to the incorporation of oxygen in both the gas and solid phases (He et al., 2017). Similarly, the addition of CO2 leads to more oxidized solid compounds (Trainer et al., 2004). Since CO2 is a more important source of oxygen than CO, one would expect oxygen incorporation to be better. In our study, this corresponds to the observation that oxidized organic compounds are better produced in the gas phase with CO2 than with CO when combined with CH4 and therefore when organic growth is not too limited. For the solid phase, Drant et al. (2025) observed in N2/CH4-rich exoplanetary haze analogs that the characteristic absorption bands of oxygenated functions (O–H, C=O) are clearly visible in analogs produced with CO2, but not or only very weakly in those produced with CO alone.
@@ 4.2 Photochemistry of hydrocarbons in temperate methane-rich exoplanet atmosphere
Experiments conducted in reduced gas mixtures reveal similar formation rates for C2 compounds, preventing any definitive conclusions about selectivity. Simulated abundance using the 0D model also predicts comparable values for these molecules, but with a slight selectivity observed for C2H6 and C2H4 relative to C2H2. This contrasts with the hydrocarbon photochemistry observed in the gas giants Jupiter and Saturn, where detected abundance is depleted in C2H4 and enriched in C2H6 and C2H2 (Moses et al., 2000, 2005; Nixon et al., 2010). These planets have hydrogen-dominated atmospheres with significant helium concentrations and a small fraction of methane (around 0.1%). They are cooler — around 150 K at 1 mbar.
The low abundance of C2H4 is explained by its high photodissociation rate at Lyman-$\alpha$, producing C2H2, and by its reaction with H, which recycles it into CH4 (Gladstone et al., 1996). In contrast, C2H6 is shielded by methane and thus remains stable at Lyman-$\alpha$. In our study, the energy distribution of electrons in our plasma covers a much wider range than that of photons in planetary atmospheres, which are largely concentrated around Lyman-$\alpha$. Consequently, the wavelength selectivity of photodissociation is not reproduced, which likely contributes to the differences observed in the relative abundance of C2 compounds between our experiments and planetary atmospheres.
Additionally, C2H2, as a highly unsaturated compound, would normally be unstable in an H2-dominated environment. However, it is efficiently recycled via the recombination reaction
$C_{2}H+H_{2}\longrightarrow C_{2}H_{2}+H.$ (5)
We noticed that this reaction is negligible in our reduced gas mixtures, whereas it becomes significant in oxidized mixtures. Since CO and CO2 are sources of carbon but create a less reducing environment than CH4, this increases the degree of unsaturation of the products. This has also been demonstrated for the production of solid products (He et al., 2017; Yang et al., 2025). This corresponds to the fact that we observe a depletion of C2H2 in mixtures containing only CH4, whereas there is no apparent selectivity in mixtures containing CO and CO2.
Temperature differences may also play a role in C2 relative abundance. Adams et al. (2022) suggests that warmer H2-rich atmospheres, compared to Titan, favor dynamic chemical cycles involving C2H2, C2H3, and C2H4 through reactions with atomic hydrogen. These processes can reduce C2H6 and C2H2 concentrations while increasing the relative abundance of C2H4.
@@ 4.3.1 Hydrocarbons
The C2 compounds appear to be relatively abundant and efficiently produced in our experiments, which are dominated by out-of-equilibrium chemistry. Similarly, a 1D non-equilibrium chemical model fitting the JWST observational data of K2-18 b predicts that C2H2, C2H4 and C2H6 could reach volume mixing ratios of up to a few percent in the upper atmosphere (at 1 mbar) (Jaziri et al., 2025). C2H2 and C2H4 are in principle detectable with JWST, using combined observations from the NIRSpec/G395 and MIRI LRS instruments, provided their mixing ratios exceed 10-6 (Gasman et al., 2022). This suggests that the detection of hydrocarbons in the atmosphere of K2-18 b is plausible if methane is abundant and predominant over carbon dioxide, as predicted by Wogan et al. (2024); Schmidt et al. (2025); Jaziri et al. (2025); Hu et al. (2025), and if photochemistry is sufficiently active. The 3.4 and 6.9 µm features observed in the NIRSpec and MIRI data could plausibly be attributed to C2H6 (Luque et al., 2025), providing an alternative to their attribution to DMS (dimethyl sulfide) or DMDS (dimethyl disulfide) by Madhusudhan et al. (2025). In summary, two-atom hydrocarbons are good candidates for detection with current JWST observations and future high-resolution ground-based observations, both in terms of expected abundance and photochemical plausibility. It is necessary to consider them before considering more complex species with a variety of potential sources.
@@ 4.3.2 Oxidized organics
Our complementary experimental approach combining MS and IR spectroscopy enables robust identification of H2CO, CH3OH, and CH3CHO formation in oxidized gas mixtures. Jaziri et al. (2025) predict that H2CO, CH3OH, and CH3CHO could reach volume mixing ratios of up to a few parts per million (10-6) in the upper atmosphere of K2-18 b. However, detectability studies with JWST indicate that species such as H2CO and CH3OH require volume mixing ratios exceeding 10-5 to produce measurable features in H2-dominated atmospheres (Huang et al., 2022; Zhan et al., 2022). Their detection in the atmosphere of K2-18 b therefore seems unlikely given current sensitivity limits. Nevertheless, our results suggest that in other H2-rich exoplanets with a higher CO and/or CO2 content, the formation of oxidized organic species may be particularly efficient. In such environments, these molecules could form more abundantly and potentially reach detectable levels with JWST or with future high-resolution ground-based instruments on the Extremely Large Telescope (ELT), such as ANDES (ArmazoNes high Dispersion Echelle Spectrograph, Palle et al. 2025) and PCS (Planetary Camera and Spectrograph, Kasper et al. 2021).
@@ 5 Conclusion
We studied out-of-equilibrium chemistry in H2-dominant atmospheres typical of temperate sub-Neptunes by combining experimental and numerical simulations. This approach complements previous laboratory studies by enabling more accurate molecular identification and clearer elucidation of reaction pathways. Our experiments, which are based on the combined use of MS and IR spectroscopy, enabled the reliable detection of reduced (C2H2, C2H4, C2H6) and oxidized (CO, CO2, H2O, H2CO, CH3OH, CH3CHO) species in gas mixtures covering a wide range of C/O ratios and metallicities. The use of the 0D photochemical model allowed us to extract abundance trends and identify formation pathways.
Hydrocarbon compounds are produced in methane-rich atmospheres in particular. The efficiency of this organic growth increases with metallicity in CH4-rich and CO-rich atmospheric analogs. In CO2-rich atmospheric analogs, oxidative loss reactions become predominant and limit the formation of hydrocarbons. The formation of hydrocarbons from methane can be explained by reaction pathways that have already been well studied in the atmospheres of the Solar System, whereas their formation from oxidized reactants — CO or CO2 — involves more reaction intermediates and a change in the predominance of certain pathways. This explains the lower abundance of reduced organic compounds in an oxidized atmosphere, combined with the emergence of oxidation pathways that consume reaction intermediates, which have more influence for low C/O ratios. The addition of oxygen diversifies the chemistry, and in CO-rich and CO2-rich mixtures in particular, it leads to the formation of the oxidized organic compounds presented above. The combination of CH4 and CO has a twofold effect. On the one hand, it provides carbon-rich atmospheres, which promote the efficient formation of hydrocarbons. On the other hand, it leads to moderately oxygenated atmospheres, which promote the formation of oxidized organic compounds without overly enhancing the pathways for hydrocarbon destruction.
Our study highlights the importance of adopting a synergistic approach combining experiments, modeling, and observations in order to deepen our understanding of exoplanet atmospheres. Each field offers unique perspectives, but through close collaboration their potential is greatly enhanced, and therefore strengthening these interdisciplinary efforts is essential.